# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

### v1

@@ 3.4.1. Blind Signatures
Blind signatures (Chaum, 1983) are a form of digital signature in which the signer produces a valid signature on a message without learning its content: before signing, the message is blinded, i.e., transformed through a random element that makes it unrecognizable. Once the contents of the signed message are unblinded, they can be publicly verified as usual.
3.4.2. Randomizable Signatures
@@
Randomized signatures provide a way to randomize a previously generated signature, transforming it into a new valid signature, but making it impossible to link the two instances together. This enables a level of anonymity by unlinking user identities from their transaction history, as each transaction appears distinct even if performed by the same user. The original idea was proposed by Camenisch and Lysyanskaya (Camenisch and Lysyanskaya, 2003, 2004) but suffered from a linear size in the number of messages to be signed. Later, Pointcheval-Sanders proposed more efficient signature and verification algorithms (Pointcheval and Sanders, 2016). Existing works (e.g., (Baudet et al., 2023; Wüst et al., 2019; Sarencheh et al., 2025; Tomescu et al., 2022)) adopt either Pointcheval-Sanders (Pointcheval and Sanders, 2016) or Coconut (Sonnino et al., 2020), a threshold scheme derived from it. The Pointcheval-Sanders scheme has the ability to produce signatures of a message directly from its commitment. This effectively makes it a blind signature, thanks to the hiding property of commitments.
3.4.3. Group Signatures
@@
Group signature schemes (Chaum and van Heyst, 1991) enable any member of a designated group to generate a valid signature without revealing who signed (ensuring anonymity). A verifier uses a group public key to confirm that a valid signature was produced by some group member. This property also implies unlinkability: it is not possible to determine whether two different signatures were produced by the same user (Perera et al., 2022). However, group signatures can be made linkable to prevent double spending in transactions (e.g., (Zhang et al., 2019)). Group signatures also support user traceability: a designated group manager holds a group secret key that enables him to revoke a signer’s anonymity and reveal his identity when necessary.
3.4.4. Ring Signatures
@@
Ring signatures (Rivest et al., 2001) are similar to group signatures, as they hide the signer’s identity behind a group. Differently from group signatures, user anonymity is permanent, because there is no revoke mechanism. Indeed, the signer simply joins their public key with a set of other public keys during the signing phase. A verifier needs the whole list of public keys to verify the validity of the signature, and cannot discern which one of the participants corresponding to the public keys is the signer, but can only guess with probability . There are some variant ring signatures, where anonymity is somewhat limited and the signer can be traced under special circumstances, e.g., after double spending. These are called traceable ring signatures (Fujisaki and Suzuki, 2007; Fujisaki, 2011). Ring signatures can also be linkable (Liu and Wong, 2005), if they allow anyone to determine if two signatures were produced by the same signer (e.g., (Van Saberhagen, 2013; Li et al., 2021)).
3.4.5. Threshold Signatures
@@
Threshold signatures are a type of signature that can be jointly produced by a committee of signers. A -of- signature scheme requires at least signers out of a total of participants to produce a valid signature. Works that use threshold signatures (Androulaki et al., 2020; Sarencheh et al., 2025; Kiayias et al., 2022; Baudet et al., 2023) use threshold blind signatures, based on Coconut (Sonnino et al., 2020).
@@ 3.6. One-time Public Keys
3.6.1. Stealth Addresses
A stealth address hides the recipient’s identity by making each payment unlinkable. It is a one-time public key generated by the transaction sender with public key provided by the transaction receiver. Recipient does not update his public key, but the sender can derive fresh one-time keys using stealth addresses. Let be a generator of an elliptic curve group with scalar field . Each receiver has a public view key and a public spend key , where are secret scalars. A sender chooses a random scalar and computes the shared secret , with a hash-to-scalar function, and the one-time public key . Then, the sender includes and in the transaction. The recipient scans the blockchain and for each and computes . If then the output belongs to the recipient.
3.6.2. Updatable Public Keys
Updatable public keys allow to update the recipient’s public key in a decentralized manner, producing dynamic and evolving public keys. Quisquis (Fauzi et al., 2019) introduces this technique to update both sender and recipient public keys. Formally, given a prime-order group with the prime order of the group, and its generator, each user has a secret key and a public key , for some random value . To perform a transaction, the sender updates the public key by choosing a random value and computing . The updated key is indistinguishable from a fresh key, thanks to the Decisional Diffie-Hellman assumption. The recipient verifies whether a given updated public key corresponds to his secret key . This works because updated keys still verify correctly with the original secret key.
4. Design Dimensions of Privacy-Preserving Digital Currencies
Dimension of analysis along which we review privacy-preserving digital currencies.
The research community has explored various solutions to ensure anonymity, confidentiality, and auditability in digital currencies. We examine the most relevant contributions from the literature, analyzing how they address key design challenges. To provide a comprehensive overview of the diverse approaches studied so far, we identify key dimensions of analysis—protocol, value representation, privacy features, and auditability—and highlight the most significant design choices. These interconnected dimensions embed privacy features into the system’s design, linking them to the ledger structure, value representation, fund exchange, and auditability. Figure 1 outlines the dimensions that structure this section. Table LABEL:tab:sota-cryptocurrencies summarizes key contributions, whereas Table 1 provides a description of the labels used in our analysis.
| Dimension | Legend | |||
| Protocol | ||||
| Type | CD | CBDC | D | Digital currency |
| Blockchain Compatibility | B | Bitcoin | E | Ethereum |
| Value Representation | ||||
| Data Model | A | Account-based | CT | Chaumian token |
| T2B | Transaction per bank table | UT | Unspent token | |
| Other | ||||
| Auditability | AB | Anonymity budget | L | Operating limits |
| RA | Revocable anonymity | TR | Coin tracing | |
| ZD | Verification without disclosure | |||
| Cryptographic Primitives | BS | Blind signature | C | Commitment |
| DS | Digital signature | GS | Group signatures | |
| HC | Homomorphic commitment | IBE | Identity-based encryption | |
| MS | Multi signature | MT | Merkle tree | |
| O-RAM | Oblivious RAM | PE | Public key encryption | |
| RS | Ring signature | SA | Stealth address | |
| UP | Updatable public keys | ZK | Zero-knowledge proof | |
| Threshold-based version of | Randomizable version of | |||
| * | Custom variant of | |||
| Generics | ✓ | Yes | No | |
| ◐ | Partial | Alternative | ||
| ○ | Abstracted | |||
@@ 4.1. Protocol
Over time, different generations of approaches have emerged, each addressing specific features and evolving needs. These developments have been driven by key factors, including the introduction of blockchain technology, advancements in cryptographic primitives, and the growing interest in designing CBDCs. The protocol dimension provides a broad perspective on the diverse range of existing solutions.
4.1.1. Generations of Privacy-preserving Digital Currencies
The term digital currency encompasses a range of different forms of money that is managed, stored, and exchanged on digital computer systems over the Internet222https://en.wikipedia.org/wiki/Digital_currency. Cryptocurrency and CBDC represent two specific forms of digital currency. A cryptocurrency is defined, stored, and exchanged over peer-to-peer systems, such as the blockchains. A CBDC is digital currency issued by a Central Bank (Bech and Garratt, 2017), where the term is often used to indicate that the currency generally does not have a classical physical form of a fiat currency. We identify three main generations of solutions for privacy-preserving digital currencies.
@@ 4.1.1. Generations of Privacy-preserving Digital Currencies
The first generation focuses on the design of digital currencies leveraging digital signatures and a centralized authority (i.e., the issuing bank) (Chaum, 1983). The fundamental problem relates to guaranteeing anonymity of sender (i.e., the payer) although he must interact with the central bank to obtain valid tokens (representing coins) to perform subsequent payments. Chaum (Chaum, 1983) proposes one of the first attempts to construct a digital currency, commonly referred to as ecash. Later, several works improved ecash with different features, including support for different coin denominations (e.g., (Dold, 2019)), partial spending (e.g., (Dold, 2019)), token transferability (e.g., (Hayes, 1990; Sai Anand and Madhavan, 2000; Baldimtsi et al., 2015; Bauer et al., 2021; Chase et al., 2014)), offline spending (e.g., (Chaum et al., 1990)), and auditability (e.g., (Camenisch et al., 2006; Dold, 2019)).
@@
The second generation of works emerged with the introduction of Bitcoin (Nakamoto, 2008). Bitcoin and, later, Ethereum (Buterin et al., 2013) propose public and permissionless blockchains, meaning that the ledger contains transactions visible to everyone and anyone can participate in the process of updating the public ledger by following a specific (validation and consensus) protocol. Being public, these blockchains use addresses (acting as pseudonyms) to exchange money in a pseudo-anonymous manner, as addresses are not directly tied to user identities. Public blockchains cannot easily and fully satisfy two privacy-related features: untraceability and unlinkability (Rahalkar and Virgaonkar, 2024). Untraceability ensures that, for any given incoming transaction, all potential senders are equally likely. Unlinkability means that, for any two outgoing transactions, there is no way to prove they were sent to the same recipient. While generating a new address for each transaction can enhance these properties, various studies indicate that achieving true anonymity and unlinkability on public blockchains remains challenging. An analysis of the transaction graph can reveal clusters of pseudonyms or patterns that link with high probability two or more addresses (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). The second generation of solutions mainly focuses on improving unlinkability, anonymity, and confidentiality, often considering decentralized architectures (e.g., (Miers et al., 2013; Van Saberhagen, 2013; Fauzi et al., 2019; Baudet et al., 2023; Friolo et al., 2024; Sarencheh et al., 2025; Androulaki et al., 2013; Herrera-Joancomartí, 2015; Meiklejohn et al., 2013)).
@@
We classify this second generation of approaches into three main categories: privacy-native protocols, mixing protocols, and off-ledger solutions. Changing the protocol of existing blockchains is hard, as it involves forks. Therefore, privacy-native protocols define (mostly) novel currencies with improved confidentiality, anonymity, and unlinkability features (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Ben Sasson et al., 2014; Tomescu et al., 2022)). Conversely, mixing protocols propose supplementary services for permissionless blockchains aimed to collect a group of transactions, scramble senders and receivers, and publish them as a single group transaction (e.g., (Maxwell, 2013; Bissias et al., 2014; Bonneau et al., 2014; Valenta and Rowan, 2015; Ruffing and Moreno-Sanchez, 2017; Ziegeldorf et al., 2018; Qin et al., 2023; Wang et al., 2023)). Mixing protocols break the link between sender and receiver, thus reducing the possibility of linking together addresses of the blockchain (i.e., they enhance unlinkability of transactions). Off-ledger solutions are a specific approach related to blockchains (which serve as Layer 1) and the ability to build decentralized peer-to-peer Layer 2 systems on top of them. Layer 2 solutions process transactions off the main chain, reducing the need to record every transaction on the shared ledger (e.g., (Qin et al., 2023; Wang et al., 2023; Baudet et al., 2023; Seres et al., 2019)). Since we focus on on-ledger privacy, we omit a detailed analysis of Layer 2 privacy solutions.
@@
While protecting the privacy and anonymity of payers and payees remains essential, the third generation of privacy-preserving digital currencies introduces mechanisms that balance these guarantees with auditability and regulatory compliance (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Narula et al., 2018; Li et al., 2021; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Yu et al., 2024; Xue et al., 2023)). These approaches are particularly motivated by concerns around AML and CFT. Auditability refers to the system’s ability to track, verify, or selectively disclose information about transactions, without fully compromising user privacy. This capability enables the enforcement of specific policies that apply to users, coins, or transactions behaviors over time. To achieve this, modern privacy-preserving systems integrate mechanisms such as revocable anonymity, trapdoor-based decryption, and ZKPs for correctness and compliance.
@@ 4.1.2. System Architecture
We examine the system architectures underlying privacy-preserving digital currencies and mixing protocols, which enhance transaction unlinkability.
@@
The first generation of privacy preserving currencies typically rely on centralized architectures involving up to three or four trusted authorities, each responsible for distinct functions. They include: (i) a central bank, which issues digital coins and maintains monetary control; (ii) a tracing authority, responsible for auditing and tracing under specific conditions; (iii) a registration authority, in charge of user on-boarding and identity verification; and, in some cases, (iv) a technical support entity, which collaborates with the tracing authority to enable conditional de-anonymization. Most early proposals assume a single central bank (e.g., (Chaum, 1983; Dold, 2019; Cecchetti et al., 2017; Zhang et al., 2019; Beer et al., 2024)). For privacy, special cryptographic primitives are used, such as blind signatures (Chaum, 1983), which allow the bank to issue valid coins without learning user identities. To support auditability and regulatory compliance, systems often include a separate tracing authority that can decrypt transactions or enforce specific policies (e.g., check or bypass an anonymity budget). This entity operates independently from the central bank, thereby preserving some separation of duties (e.g., (Sander and Ta-Shma, 1999; Wüst et al., 2022)). Yu et al. (Yu et al., 2024) introduce a technical support authority that must cooperate with the tracing authority to extract sender identities. In the context of CBDCs, system architectures usually include a registration authority, responsible for enrolling user registration and managing identity records (e.g., (Zhang et al., 2019))—an essential requirement considering regulatory frameworks (e.g., AML and CFT).
@@
Blockchains introduced a fundamental shift in the design of digital currencies by leveraging decentralization and cryptographic protocols. Decentralization serves two main purposes. First, it enhances resilience against malicious actors or censorship by distributing trust across multiple entities. This is often supported by cryptographic techniques such as threshold encryption, secret sharing, and multi-signature schemes. Second, it enables intermediary-free value exchange, using commitments and digital signatures to ensure integrity, authenticity, and non-repudiation. Many privacy-native solutions adopt decentralized architectures in line with these principles (e.g., (Noether et al., 2016; Van Saberhagen, 2013; Poelstra, 2016; Androulaki et al., 2020; Narula et al., 2018; Ma et al., 2021; Xue et al., 2023)). Moreover, the programmability of blockchains has enabled privacy-preserving overlays (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)) and smart contracts333Although smart contracts are logically centralized, they are executed on each node; therefore, we consider smart contracts as a decentralized solution. implementing private tokens “pegged” to the coin of the primary chain (e.g., (Bünz et al., 2020; Baudet et al., 2023; Sarencheh et al., 2025)).
@@
Different architectural styles have been explored for mixing protocols, including both centralized (e.g., (Bonneau et al., 2014; Valenta and Rowan, 2015; Heilman et al., 2016; Qin et al., 2023; Liu et al., 2018)) and decentralized solutions (e.g., (Ruffing et al., 2014; Ziegeldorf et al., 2018; Bissias et al., 2014; Ruffing and Moreno-Sanchez, 2017; Nosouhi et al., 2023; Wang et al., 2023)). In centralized solutions, a third party (often called mixer or tumbler) provides the mixing service for a fee. While such designs are simple and support large anonymity sets, they introduce risks such as availability, theft, and potential privacy breaches. Accordingly, different trust models have been proposed to mitigate these concerns (see Section 4.1.4). Decentralized mixers remove the need for a trusted party by using cryptographic protocols or peer-to-peer coordination, enhancing security and censorship resistance, but they often incur higher complexity, require user coordination, and typically support smaller anonymity sets. Notable examples include Xim (Bissias et al., 2014), CoinShuffle++ (Ruffing et al., 2016), and ValueShuffle (Ruffing and Moreno-Sanchez, 2017). Both centralized and decentralized approaches face challenges such as Sybil attacks and regulatory scrutiny. In a Sybil attack, an adversary inject fake participants into the mix to degrade anonymity.

### v2

@@ 3.4. Anonymity-enhanced Signatures
(Chaum, 1983) are a form of digital signature in which the signer produces a valid signature on a message without learning its content: before signing, the message is blinded, i.e., transformed through a random element that makes it unrecognizable. Once the contents of the signed message are unblinded, they can be publicly verified as usual.
- Randomizable signatures:
@@
provide a way to randomize a previously generated signature, transforming it into a new valid signature, but making it impossible to link the two instances together. This enables a level of anonymity by unlinking user identities from their transaction history, as each transaction appears distinct even if performed by the same user. The original idea was proposed by Camenisch and Lysyanskaya (Camenisch and Lysyanskaya, 2003, 2004). Later, Pointcheval-Sanders proposed more efficient signature and verification algorithms (Pointcheval and Sanders, 2016). Existing works (e.g., (Baudet et al., 2023; Wüst et al., 2019; Sarencheh et al., 2025; Tomescu et al., 2022)) adopt either Pointcheval-Sanders or Coconut (Sonnino et al., 2020), a threshold scheme derived from it. The Pointcheval-Sanders scheme can produce signatures of a message directly from its commitment; this effectively makes it a blind signature, thanks to the hiding property of commitments.
- Group signatures:
@@
schemes (Chaum and van Heyst, 1991) enable any member of a designated group to generate a valid signature without revealing who signed (ensuring anonymity). A verifier uses a group public key to confirm that a valid signature was produced by some group member. This property also implies unlinkability: it is not possible to determine whether two different signatures were produced by the same user (Perera et al., 2022). However, group signatures can be made linkable to prevent double spending in transactions (e.g., (Zhang et al., 2019)). Group signatures also support user traceability: a designated group manager holds a group secret key that enables him to revoke a signer’s anonymity and reveal his identity when necessary.
- Ring signatures:
@@
(Rivest et al., 2001) are similar to group signatures, as they hide the signer’s identity behind a group. Differently from group signatures, user anonymity is permanent, because there is no revoke mechanism. Indeed, the signer simply joins their public key with a set of other public keys during the signing phase. A verifier needs the whole list of public keys to verify the signature validity, and cannot discern which public key is of the signer, but can only guess . There are some variant ring signatures, where anonymity is somewhat limited and the signer can be traced under special circumstances, e.g., after double spending. These are called traceable ring signatures (Fujisaki and Suzuki, 2007; Fujisaki, 2011). Ring signatures can also be linkable (Liu and Wong, 2005), if they allow anyone to determine if two signatures were produced by the same signer (e.g., (Van Saberhagen, 2013; Li et al., 2021)).
- Threshold signatures:
@@
are a type of signature that can be jointly produced by a committee of signers. A -of- signature scheme requires at least signers out of a total of participants to produce a valid signature. Works that use threshold signatures (Androulaki et al., 2020; Sarencheh et al., 2025; Kiayias et al., 2022; Baudet et al., 2023) use threshold blind signatures, based on Coconut (Sonnino et al., 2020).
@@ 3.6. One-time Public Keys
- Stealth addresses:
hide the recipient’s identity making each payment unlinkable. They are one-time public keys generated by the transaction sender with public key provided by the receiver. Recipient does not update his public key, but the sender can derive fresh one-time keys for each transaction. Let be a generator of an elliptic curve group with scalar field . Each receiver has a public view key and a public spend key , where are secret scalars. A sender chooses a random scalar and computes the shared secret , with a hash-to-scalar function, and the one-time public key . The sender includes and in the transaction. The recipient can scan the ledger and for each and computes . If then the output belongs to the recipient.
- Updatable public keys:
allow to update the recipient’s public key in a decentralized manner. Quisquis (Fauzi et al., 2019) introduces this technique to update both sender and recipient public keys. Given a prime-order group with the prime order of the group, and its generator, each user has a secret key and a public key , for some random value . To perform a transaction, the sender updates the public key by choosing a random value and computing . The updated key is indistinguishable from a fresh key. The recipient verifies whether a given updated public key corresponds to his secret key . This works because updated keys still verify correctly with the original secret key.
4. Design Dimensions of Privacy-Preserving Digital Payment Systems
The research community has explored various solutions to ensure anonymity, confidentiality, and auditability in crypto-assets and digital payment systems. We identify key dimensions of analysis—protocol, value representation, privacy features, and auditability—and highlight the most significant design choices. These dimensions are interconnected and jointly embed privacy features into the system. Figure 1 outlines the dimensions that structure this section. Table 1 describes the labels used in Table LABEL:tab:sota-cryptocurrencies, which summarizes our literature review.
@@ 4.1. Protocol
The protocol dimension outlines the main system design choices through which existing solutions can be compared.
Dimension Legend Protocol Type CD CBDC D Digital payment system Blockchain Compatibility B Bitcoin E Ethereum Value Representation Data Model A Account-based CT Chaumian token T2B Transaction per bank table UT Unspent token Other Auditability AB Anonymity budget L Operating limits RA Revocable anonymity TR Coin tracing ZD Verification without disclosure Cryptographic Primitives BS Blind signature C Commitment DS Digital signature GS Group signatures HC Homomorphic commitment IBE Identity-based encryption MS Multi signature MT Merkle tree O-RAM Oblivious RAM PE Public key encryption RS Ring signature SA Stealth address UP Updatable public keys ZK Zero-knowledge proof Threshold-based version of Randomizable version of * Custom variant of Generics ✓ Yes No ◐ Partial Alternative ○ Abstracted
4.1.1. Generations of Privacy-preserving Digital Payment Systems
Privacy-preserving digital payment systems have evolved through three generations, shaped by blockchain adoption, advances in cryptography, and interest in CBDCs. The first generation focuses on centralized authorities, emphasizing sender anonymity. The second, emerging with Bitcoin, enables decentralized value transfer and focuses on unlinkability and untraceability. The third adds mechanisms to balance privacy with auditability and regulatory compliance.
@@ 4.1.1. Generations of Privacy-preserving Digital Payment Systems
The first generation focuses on the design of privacy-preserving value transfer systems relying on digital signatures and a centralized authority (i.e., the issuing bank) (Chaum, 1983). The fundamental problem relates to guaranteeing anonymity of sender (i.e., the payer) although he must interact with the central bank to obtain valid tokens (representing coins) to perform subsequent payments. Chaum (Chaum, 1983) introduces one of the first untraceable payment systems, commonly known as ecash. Later, several works improved ecash with different features (Andola et al., 2021), including support for different coin denominations (e.g., (Dold, 2019)), partial spending (e.g., (Dold, 2019)), token transferability (e.g., (Hayes, 1990; Sai Anand and Madhavan, 2000; Baldimtsi et al., 2015; Bauer et al., 2021; Chase et al., 2014)), offline spending (e.g., (Chaum et al., 1990)), and auditability (e.g., (Camenisch et al., 2006; Dold, 2019)).
@@
The second generation emerged with the introduction of Bitcoin (Nakamoto, 2008). Bitcoin and, later, Ethereum (Buterin et al., 2013) propose public and permissionless blockchains, meaning that the ledger contains transactions visible to everyone and anyone can participate in the process of updating the public ledger. Being public, these blockchains use addresses to exchange assets in a pseudo-anonymous manner, as addresses are not directly tied to user identities. Public blockchains cannot easily and fully satisfy two privacy-related features: untraceability and unlinkability (Rahalkar and Virgaonkar, 2024). Untraceability ensures that, for any given incoming transaction, all potential senders are equally likely. Unlinkability means that, for any two outgoing transactions, there is no way to prove they were sent to the same recipient. While generating a new address for each transaction can enhance these properties, various studies indicate that achieving true anonymity and unlinkability on public blockchains remains challenging. An analysis of the transaction graph can reveal clusters of pseudonyms or patterns that link with high probability two or more addresses (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). The second generation of solutions mainly focuses on improving unlinkability, anonymity, and confidentiality, often considering decentralized architectures (e.g., (Miers et al., 2013; Van Saberhagen, 2013; Fauzi et al., 2019; Baudet et al., 2023; Friolo et al., 2024; Sarencheh et al., 2025; Androulaki et al., 2013; Herrera-Joancomartí, 2015; Meiklejohn et al., 2013)).
@@
We classify this second generation of approaches into three main categories: privacy-native protocols, mixing protocols, and off-ledger solutions. Changing the protocol of existing blockchains is hard, as it involves forks. Therefore, privacy-native protocols (mostly) defines novel solutions with improved confidentiality, anonymity, and unlinkability features (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Ben Sasson et al., 2014; Tomescu et al., 2022)). Conversely, mixing protocols propose supplementary services for permissionless blockchains aimed to collect a group of transactions, scramble senders and receivers, and publish them as a single group transaction (e.g., (Maxwell, 2013; Bissias et al., 2014; Bonneau et al., 2014; Valenta and Rowan, 2015; Ruffing and Moreno-Sanchez, 2017; Ziegeldorf et al., 2018; Qin et al., 2023; Wang et al., 2023)). Mixing protocols break the link between sender and receiver, thus reducing the possibility of linking together addresses of the blockchain. Off-ledger solutions are a specific approach related to blockchains (that serve as Layer 1) and the ability to build decentralized peer-to-peer Layer 2 systems on top of them. Layer 2 solutions process transactions off the main chain, reducing the need to record every transaction on the shared ledger (e.g., (Qin et al., 2023; Wang et al., 2023; Baudet et al., 2023; Seres et al., 2019)). Since we focus on on-ledger privacy, we omit a detailed analysis of Layer 2 privacy solutions.
@@
While protecting the privacy and anonymity of payers and payees remains essential, the third generation of solutions introduces mechanisms that balance these guarantees with auditability and regulatory compliance (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Narula et al., 2018; Li et al., 2021; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Yu et al., 2024; Xue et al., 2023)). These approaches are motivated by concerns around AML and CFT. Auditability refers to the system’s ability to track, verify, or selectively disclose information about transactions, without fully compromising user privacy. This enables the enforcement of specific policies that apply to users, coins, or transactions behaviors over time. To this end, modern privacy-preserving systems integrate mechanisms such as revocable anonymity, trapdoors, and ZKPs for correctness and compliance.
@@ 4.1.2. System Architecture
We examine the architecture of privacy-preserving payment systems and protocols.
@@
The first generation of privacy preserving solutions typically rely on centralized architectures involving up to four trusted authorities, each responsible for distinct functions. They include: (i) a central bank, which issues digital coins and maintains monetary control; (ii) a tracing authority, responsible for auditing and tracing under specific conditions; (iii) a registration authority, in charge of user on-boarding and identity verification; and, in some cases, (iv) a technical support entity, which collaborates with the tracing authority to enable conditional de-anonymization. Most early proposals assume a single central bank (e.g., (Chaum, 1983; Dold, 2019; Cecchetti et al., 2017; Zhang et al., 2019; Beer et al., 2024)). For privacy, special cryptographic primitives are used, such as blind signatures (Chaum, 1983), which allow the bank to issue valid coins without learning user identities. To support auditability and regulatory compliance, systems often include a separate tracing authority that can decrypt transactions or enforce specific policies (e.g., check or bypass an anonymity budget). This entity operates independently from the central bank, thereby preserving some separation of duties (e.g., (Sander and Ta-Shma, 1999; Wüst et al., 2022)). Yu et al. (Yu et al., 2024) introduce a technical support authority that must cooperate with the tracing authority to extract sender identities. In the context of CBDCs, system architectures usually include a registration authority, responsible for enrolling user registration and managing identity records (e.g., (Zhang et al., 2019))—an essential requirement considering regulatory frameworks (e.g., AML and CFT).
@@
Blockchains introduced a fundamental shift in system design by leveraging decentralization, which serves two main purposes. First, it enhances resilience against malicious actors or censorship by distributing trust across multiple entities. Second, it enables intermediary-free value exchange, using commitments and digital signatures to ensure integrity, authenticity, and non-repudiation. Many privacy-native solutions adopt decentralized architectures in line with these principles (e.g., (Noether et al., 2016; Van Saberhagen, 2013; Poelstra, 2016; Androulaki et al., 2020; Narula et al., 2018; Ma et al., 2021; Xue et al., 2023)). Furthermore, the programmability of blockchains has enabled privacy-preserving overlays (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)) and smart contracts333Although logically centralized, smart contracts are executed on each node; therefore, we consider them as a decentralized solution. implementing private tokens “pegged” to the coin of the primary chain (e.g., (Bünz et al., 2020; Baudet et al., 2023; Sarencheh et al., 2025)).
@@
Both centralized (e.g., (Bonneau et al., 2014; Valenta and Rowan, 2015; Heilman et al., 2016; Qin et al., 2023; Liu et al., 2018)) and decentralized solutions (e.g., (Ruffing et al., 2014; Ziegeldorf et al., 2018; Bissias et al., 2014; Ruffing and Moreno-Sanchez, 2017; Nosouhi et al., 2023; Wang et al., 2023)) have been proposed to implement mixing protocols. In centralized solutions, a third party (mixer or tumbler) provides the mixing service for a fee. While such designs are simple and support large anonymity sets, they introduce risks such as availability, theft, and potential privacy breaches. Therefore, different trust models have been proposed to mitigate these concerns (see Section 4.1.4). Decentralized mixers remove the need for a trusted party by using cryptographic protocols or peer-to-peer coordination, enhancing security and censorship resistance, but they often incur higher complexity, require user coordination, and typically support smaller anonymity sets (e.g., Xim (Bissias et al., 2014), CoinShuffle++ (Ruffing et al., 2016), and ValueShuffle (Ruffing and Moreno-Sanchez, 2017)). Both centralized and decentralized approaches face challenges such as Sybil attacks444In a Sybil attack, an adversary injects fake participants into the mix to degrade anonymity. and regulatory scrutiny.